3. At points A, B and C of a rod, which are in a straight line, the parallel forces F1 = 40N, F2 = 30N and F3 = 30N operate in the same directions. Find the resultant of the three forces, knowing that AB = 28cm and AC = 60cm. Point B is located between A and C.


1
Expert's answer
2021-03-24T19:42:33-0400

Since the forces are parallel and have the same direction, then


"F_{res}=F_1+F_2+F_3=40+30+30=100\ (N)" . Answer